OpenStudy (anonymous):

simplify 3^1/2 . 3^1/2

OpenStudy (mertsj):

Use the principle: ( a^m)(a^n)=a^(m+n)

OpenStudy (mertsj):

\[a^ma^n=a ^{m+n}\]

OpenStudy (anonymous):

how do you do that?

OpenStudy (mertsj):

If the bases are the same, you multiply expressions by adding the exponents. Your bases are the same. Add the exponents.

OpenStudy (anonymous):

9^1/2?

OpenStudy (mertsj):

what is the exponent of 3^1/2?

OpenStudy (anonymous):

6?

OpenStudy (mertsj):

\[4^7\]

OpenStudy (mertsj):

In the expression I just posted, 4 is the base and 7 is the exponent

OpenStudy (anonymous):

Oh okay, so then what?

OpenStudy (mertsj):

\[(4^7 )(4^6)=4^{7+6}=4^{13}\]

OpenStudy (mertsj):

Add the exponents.

OpenStudy (anonymous):

oh okay so 4^13 is the answer?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Did you see what I posted?

OpenStudy (anonymous):

yeah but is the answer 4^13?

OpenStudy (mertsj):

That is the answer to the problem I made up. Now you use the same ideas to do your problem.

OpenStudy (anonymous):

okay but how do you add fractions?

OpenStudy (mertsj):

Oh please!! Are you telling me that if you have 1/2 dollar and someone gives you 1/2 dollar you don't know how much money you now have?????

OpenStudy (anonymous):

oh so the answer to my proclem is 3^1?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

oh okay, thanks!

OpenStudy (mertsj):

yw